OpenStudy (anonymous):

Alexis calculates the surface area of a gift box as 600square inches. The actual surface area of the gift box is 592 square inches. Find the relative error of Alexis' calculation expressed to as a decimal to the nearest thousandth.

OpenStudy (anonymous):

The relative error is given by the relative difference, which is defined as\[\frac{(final)-(initial)}{(initial)}\]so here,\[\frac{600-592}{592}=0.014\]expressed to the nearest thousandth.